The beam emerging from a ruby laser (wavelength 694.3 nm) passes through a circular aperture 4.4 mm in diameter. (a) If the spread of the beam is limited only by diffraction, what is the angular spread of the beam?(b) If the beam is aimed at the Moon, how large a spot would be illuminated on the Moon's surface?
Answer #1

The beam emerging from a ruby laser having wavelegth 694.30 nm is passing through circular aperture of   diameter 4.4 mm

Diameter = 4.4 mm = 4.4*10-3 m

a)   sin\theta = 1.22\lambda /b

= 1.22*694*10-9 m/(4.4*10-3 m)

= 192.427*10-6 rad

b) The Moon is at a larger distance from a hole i.e. Earth's Surface(D>>>>b) as D is 384400 km, So Radius of image disc = (1.22*\lambdaD)/b = 1.22*3.844*108*694.3*10-9/(4.4*10-3)

= 7.4*104m